During a University of Nevada, Reno blood donation campaign, the frequencies of alleles determining the ABO blood groups were 0.70 for i, 0.2 for IA and 0.1 for IB. Assuming random mating, what is the expected frequency of blood type B?During a University of Nevada, Reno blood donation campaign, the frequencies of alleles determining the ABO blood groups were
Answer #1

Allele frequency of i (r)= 0.7

Allele frequency of IA (p) = 0.2

Allele frequency of IB (q)= 0.1

Blood type B genotypes are either IBIB or IBIi

Frequency of IBIB = q2 = 0.1*0.1= 0.01

Frequency of IBIi = 2qr = 2*0.1*0.7= 0.14

Thus expected frequency of blood type B is 0.15